Alpha Ltd is expecting an annual Earnings Before Interest and Tax of ₹1,50,000. The company had 10% debentures of ₹ 5 Lakhs. Calculate the overall cost of capital of the firm if the cost of Equity is 12% assuming NI approach in the calculation. Suggest what happens to the value of the firm when more debentures are issued.
Answer #1

Net income approach suggests that value of firm can be increased by decreasing the overall cost of capital through higher debt proportion.

EBIT 150000 Less : Interest cost 50000 (500000*10%) = 100000 i.e Earnings after tax

Market value of equity shares = Earnings after tax÷ Cost of equity. i.e 100000/12% = 833333.33

Market value of debt = 500000

Total Market value of firm = 1333333.33

Overall cost of capital in percentage = EBIT÷ Total Market value of firm =150000÷ 1333333.33 = 11.25%

When more debentures are issued the value of firm will initially increase upto a certain thereafter it will start falling.Since higher interest payment and return of shareholders ultimately affects the debt financing.